# Simulating Holographic Conformal Field Theories on Hyperbolic Lattices

### v1

@@ Abstract
We demonstrate how table-top settings combining hyperbolic lattices with nonlinear dynamics universally lead to a genuine realization of the bulk-boundary-correspondence between gravity in anti-de-Sitter (AdS) space and conformal field theory (CFT). Our concrete and broadly applicable holographic toy model simulates gravitational self-interactions in the bulk and features an emergent CFT with nontrivial correlations on the boundary. We measure the CFT data contained in the two- and three-point functions and clarify how a thermal CFT is simulated through an effective black hole geometry. As a concrete example, we propose and simulate an experimentally feasible protocol to measure the holographic CFT using electrical circuits.
@@
$=\int\frac{\mbox{d}^{2}z}{(1-|z|^{2})^{2}}\Bigl{[}\frac{1}{2}\phi(-\square+m^{2})\phi+\frac{u}{3!}\phi^{3}+\frac{\mu}{4!}\phi^{4}+\dots\Bigr{]}.$
We may interpret $\phi(z)$ as graviton-like mode, since it parametrizes small metric fluctuations. Importantly, dynamical gravity corresponds here to a real scalar field $\phi$ propagating in its own hyperbolic background with $\square=(2\ell)^{-2}(1-|z|^{2})^{2}\nabla^{2}$ the Laplace–Beltrami operator [ingredient (i)], and nonlinear terms like $u\phi^{3}$ and $\mu\phi^{4}$ correspond to gravitational self-interactions [ingredient (ii)]. We choose $m^{2}\ell^{2}>-1/4$ above the Breitenlohner–Freedman bound for a stable theory [33].
@@
$S(\{\phi_{\mu}\})=-\frac{t}{2}\sum_{\mu,\nu}\phi_{\mu}A_{\mu\nu}\phi_{\nu}+\sum_{\mu}\Bigl{(}\frac{\hat{m}^{2}}{2}\phi_{\mu}^{2}+\frac{u}{3!}\phi_{\mu}^{3}\Bigr{)}.$ (3)
Herein, $\phi_{\mu}=\phi(z_{\mu})$ is the field variable defined on the sites $z_{\mu}$ of a graph or lattice $\mathcal{G}$ with adjacency matrix $A$. Equation (3) represents a generic tight-binding Hamiltonian realizable on the platforms discussed in the introduction, and is a universal low-energy limit of more complicated Hamiltonians. The parameters $t$, $\hat{m}^{2}$, and $u$ are tunable in experiment. We neglect higher-order interaction terms beyond $\phi^{3}$ for simplicity.
The choice of graph $\mathcal{G}$ determines the curved bulk space on which the action is simulated. Two-dimensional imaginary-time AdS space is emulated by hyperbolic $\{p,q\}$ lattices with $(p-2)(q-2)>4$, labelled type-I hereafter. Black holes are realized by identifying points on two geodesics in a type-I geometry. This is equivalent to a type-II ring graph [35] obtained by a squash-and-wrap procedure, see Fig. 1. The squash step is $z_{\mu}\mapsto\zeta_{\mu}=\frac{2}{\pi}\ln(\frac{1+z_{\mu}}{1-z_{\mu}})$, which produces an infinite strip, from which a finite strip of width $w$ is obtained with its truncated edges identified, followed by the wrap step $\zeta_{\mu}\mapsto\hat{z}_{\mu}=e^{2\pi{\rm i}(\zeta_{\mu}+{\rm i})/w}$ yielding a ring. When applied to the Poincaré disk metric, this results in a timeslice of the three-dimensional Bañadas–Teitelboim–Zanelli (BTZ) black hole metric [62, 63, 64, 35, 68] given by
@@
[ingredient (iv)]. An intriguing question in holography concerns the nature of the fields $\mathcal{O}_{a}$. We leave it for future work to explore this aspect in our lattice model.
@@
Here $\theta_{a,b}$ are the angle coordinates of the boundary sites. This emulates CFTs on a circle at zero and finite temperature. While Eqs. (9) are good approximations on type-II lattices, especially for the 2-pt function, the proper quantitative formula requires to replace $\theta_{a}-\theta_{b}\to\theta_{a}-\theta_{b}+2\pi n$ in $d_{ab}$ with a subsequent sum over $n\in\mathbb{Z}$ to make the functions periodic [68]. This behavior, reminiscent of the method of images, also arises in the BTZ geometry and is thus expected here [64].
Our main novel findings from the analysis of correlations are that (a) the scaling dimension $\Delta(m^{2}\ell^{2})$ extracted from the 2-pt function also captures the scaling of the 3-pt function, indicating a consistent CFT [Fig 3d], (b) the parameter $k$ on type-II lattices determines the temperature consistent with the formula $T\ell=kP/8\pi$ [Fig. 2d], but leaves the CFT data invariant, and (c) the CFT data $\Delta(m^{2}\ell^{2})$ and $C_{3}(m^{2}\ell^{2})$ for the CFTs simulated on both type-I and type-II lattices is identical [Fig. 3d,e]. The latter finding indicates that they are the same CFTs, but at zero and finite temperature. These nontrivial results also solidify the interpretation of the type-II lattice as a geometry that emulates a black hole.
@@
with $h=0.497$ [22, 19, 68], see Fig. 2b. The value of $h$ and form of $f(X)$ depend on the $\{p,q\}$ lattice, with the universal continuum limit recovered for $h\to 0$. To extract the 3-pt coefficient $C_{3}$ in a manner that reduces scatter due to the lattice discretization, we construct the function $F_{abc}=\langle\mathcal{O}_{a}\mathcal{O}_{b}\mathcal{O}_{c}\rangle/[\langle\mathcal{O}_{a}\mathcal{O}_{b}\rangle\langle\mathcal{O}_{a}\mathcal{O}_{c}\rangle\langle\mathcal{O}_{b}\mathcal{O}_{c}\rangle]^{1/2}$ that we average over $a,b,c$ to obtain $F_{abc}\simeq C_{3}$ [68].
@@
To realize the terms in Eq. (12), we use resistors to generate the linear couplings $\sum_{\nu}(-A_{\mu\nu}+\hat{m}^{2}\delta_{\mu\nu})V_{\nu}$, while diodes can be used for the nonlinear term $uV_{\mu}^{2}$. Indeed, the current through a diode is approximated by the Shockley equation $I(V)=I_{\rm S}(\mathrm{e}^{V/V_{\rm S}}-1)\approx I_{\rm S}[V/V_{\rm S}+(V/V_{\rm S})^{2}/2]$, which yields the desired $V_{\mu}^{2}$-term after absorbing the linear part into $\hat{m}^{2}V_{\mu}$. The concrete circuit parameters are listed in Suppl. Sec. S8 [68].
@@
In Fig. 2a we present results for the so-obtained 2-pt function using an LTspice simulation of an electrical circuit with a realistic Schottky diode (model RBE1VAM20A). In Fig. 3b,c we present the 3-point function for the same parameters from a Mathematica simulation of Eqs. (12)-(14). Since the signals in actual electrical circuits are known to be close to such numerical simulations [55, 56, 57], this serves as a proof of principle that our protocol is experimentally feasible.
@@
$S_{\rm grav}[\phi]=\int\mbox{d}^{2}s\ e^{\varphi_{\star}}\Bigl{[}-\frac{1}{2}\phi(e^{-\varphi_{\star}}\nabla^{2})\phi+\frac{2}{\ell^{2}}(e^{\phi}-\phi)\Bigr{]}.$ (S6)
@@
$S_{\rm grav}[\phi]={}$ $\int\mbox{d}^{2}s\ e^{\varphi_{\star}}\Bigl{[}-\frac{1}{2}\phi(e^{-\varphi_{\star}}\nabla^{2})\phi$

### v2

@@ Abstract
We demonstrate how table-top settings combining hyperbolic lattices with nonlinear dynamics universally encode aspects of the bulk-boundary-correspondence between gravity in anti-de-Sitter (AdS) space and conformal field theory (CFT). Our concrete and broadly applicable holographic toy model simulates gravitational self-interactions in the bulk and features an emergent CFT with nontrivial correlations on the boundary. We measure the CFT data contained in the two- and three-point functions and clarify how a thermal CFT is simulated through an effective black hole geometry. As a concrete example, we propose and simulate an experimentally feasible protocol to measure the holographic CFT using electrical circuits.
@@
$=\int\frac{\mbox{d}^{2}z}{(1-|z|^{2})^{2}}\Bigl{[}\frac{1}{2}\phi(-\square+m^{2})\phi+\frac{\phi^{3}}{3\ell^{2}}+\frac{\phi^{4}}{12\ell^{2}}+\dots\Bigr{]}.$
We may interpret $\phi(z)$ as graviton-like mode, since it parametrizes small metric fluctuations. Importantly, dynamical gravity corresponds here to a real scalar field $\phi$ propagating in its own hyperbolic background with $\square=(2\ell)^{-2}(1-|z|^{2})^{2}\nabla^{2}$ the Laplace–Beltrami operator [ingredient (i)], and nonlinear terms like $\phi^{3}$ and $\phi^{4}$ correspond to gravitational self-interactions [ingredient (ii)]. We choose $m^{2}\ell^{2}>-1/4$ above the Breitenlohner–Freedman bound for a stable theory [33].
@@
$S(\{\phi_{\mu}\})=-\frac{1}{2}\sum_{\mu,\nu}\phi_{\mu}A_{\mu\nu}\phi_{\nu}+\sum_{\mu}\Bigl{(}\frac{\hat{m}^{2}}{2}\phi_{\mu}^{2}+\frac{u}{3!}\phi_{\mu}^{3}\Bigr{)}.$ (3)
Herein, $\phi_{\mu}=\phi(z_{\mu})$ is the field variable defined on the sites $z_{\mu}$ of a graph or lattice $\mathcal{G}$ with adjacency matrix $A$. Equation (3) represents a generic tight-binding Hamiltonian realizable on the platforms discussed in the introduction, and is a universal low-energy limit of more complicated Hamiltonians. The parameters $\hat{m}^{2}$ and $u$ are tunable in experiment and can be matched to the Liouville action (1) [68]. We neglect higher-order interaction terms beyond $\phi^{3}$ for simplicity, resulting in a model that deviates from Liouville gravity through this omission.
The choice of graph $\mathcal{G}$ determines the curved bulk space on which the action is simulated. Two-dimensional imaginary-time AdS space is emulated by hyperbolic $\{p,q\}$ lattices with $(p-2)(q-2)>4$, labelled type-I hereafter. Black holes are realized by identifying points on two geodesics in a type-I geometry. This is equivalent to a type-II ring graph [35] obtained by a squash-and-wrap procedure, see Fig. 1. The squash step is $z_{\mu}\mapsto\zeta_{\mu}=\frac{2}{\pi}\ln(\frac{1+z_{\mu}}{1-z_{\mu}})$, which produces an infinite strip, from which a finite strip of width $w$ is obtained with its truncated edges identified, followed by the wrap step $\zeta_{\mu}\mapsto\hat{z}_{\mu}=e^{2\pi{\rm i}(\zeta_{\mu}+{\rm i})/w}$ yielding a ring. When applied to the Poincaré disk metric, this results in a timeslice of the three-dimensional Bañados–Teitelboim–Zanelli (BTZ) black hole metric [62, 63, 64, 35, 68] given by
@@
[ingredient (iv)]. Properties like spin or scaling behavior of $\mathcal{O}_{a}$ are determined by the bulk theory, but no action underlying the CFT is specified by the duality per se. An important future research inquiry is to explore whether a representative CFT action can be constructed in the holographic toy model. Furthermore, adding more fields and symmetries to the bulk action, the consistency of richer dual CFTs can be probed.
@@
Here $\theta_{a,b}$ are the angle coordinates of the boundary sites. This emulates CFTs on a circle at zero and finite temperature. While Eqs. (9) are good approximations on type-II lattices, especially for the 2-point function, the proper quantitative formula requires to replace $\theta_{a}-\theta_{b}\to\theta_{a}-\theta_{b}+2\pi n$ in $d_{ab}$ with a subsequent sum over $n\in\mathbb{Z}$ to make the functions periodic [68]. This behavior, reminiscent of the method of images, also arises in the BTZ geometry and is thus expected here [64].
Our main novel findings from the analysis of correlations are that (a) the scaling dimension $\Delta(m^{2}\ell^{2})$ extracted from the 2-point function also captures the scaling of the 3-point function, indicating a consistent CFT [Fig 3d], (b) the parameter $k$ on type-II lattices determines the temperature consistent with the formula $T\ell=kP/8\pi$ [Fig. 2d], but leaves the CFT data invariant, and (c) the CFT data $\Delta(m^{2}\ell^{2})$ and $C_{3}(m^{2}\ell^{2})$ for the CFTs simulated on both type-I and type-II lattices is identical [Fig. 3d,e]. The latter finding indicates that they are the same CFTs, but at zero and finite temperature. These nontrivial results also solidify the interpretation of the type-II lattice as a geometry that emulates a black hole.
@@
with $h=0.497$ [22, 19, 68], see Fig. 2b. The value of $h$ and form of $f(X)$ depend on the $\{p,q\}$ lattice, with the universal continuum limit recovered for $h\to 0$. To extract the 3-point coefficient $C_{3}$ in a manner that reduces scatter due to the lattice discretization, we construct the function $F_{abc}=\langle\mathcal{O}_{a}\mathcal{O}_{b}\mathcal{O}_{c}\rangle/[\langle\mathcal{O}_{a}\mathcal{O}_{b}\rangle\langle\mathcal{O}_{a}\mathcal{O}_{c}\rangle\langle\mathcal{O}_{b}\mathcal{O}_{c}\rangle]^{1/2}$ that we average over $a,b,c$ to obtain $F_{abc}\simeq C_{3}$ [68].
@@
To realize the terms in Eq. (12), we use resistors to generate the linear couplings $\sum_{\nu}(-A_{\mu\nu}+\hat{m}^{2}\delta_{\mu\nu})V_{\nu}$, while diodes can be used for the nonlinear term $uV_{\mu}^{2}$. Indeed, the current through a diode is approximated by the Shockley equation $I(V)=I_{\rm S}(\mathrm{e}^{V/V_{\rm S}}-1)\approx I_{\rm S}[V/V_{\rm S}+(V/V_{\rm S})^{2}/2]$, which yields the desired $V_{\mu}^{2}$-term after absorbing the linear part into $\hat{m}^{2}V_{\mu}$. The concrete circuit parameters are listed in Suppl. Sec. S8 [68]. On short time scales, dissipative terms present in any realistic circuit lead to transient behavior. These effects are neglected here and the correspondence is realized in the steady state, assuming an instantaneous response to $J_{a}(t)$.
@@
In Fig. 2a we present results for the so-obtained 2-point function using an LTspice simulation of an electrical circuit with a realistic Schottky diode (model RBE1VAM20A). In Fig. 3b,c we present the 3-point function for the same parameters from a Mathematica simulation of Eqs. (12)-(14). Since the signals in actual electrical circuits are known to be close to such numerical simulations [55, 56, 57], this serves as a proof of principle that our protocol is experimentally feasible.
@@
$S_{\rm grav}[\phi]=\int\mbox{d}^{2}z\ e^{\varphi_{\star}}\Bigl{[}-\frac{1}{2}\phi(e^{-\varphi_{\star}}\nabla^{2})\phi+\frac{2}{\ell^{2}}(e^{\phi}-\phi)\Bigr{]}.$ (S6)
@@
$S_{\rm grav}[\phi]={}$ $\int\mbox{d}^{2}z\ e^{\varphi_{\star}}\Bigl{[}-\frac{1}{2}\phi(e^{-\varphi_{\star}}\nabla^{2})\phi$
@@
To connect the Liouville action to the lattice action
$S(\{\phi_{\mu}\})=-\frac{1}{2}\sum_{\mu,\nu}\phi_{\mu}A_{\mu\nu}\phi_{\nu}+\sum_{\mu}\Bigl{(}\frac{\hat{m}^{2}}{2}\phi_{\mu}^{2}+\frac{u}{3!}\phi_{\mu}^{3}\Bigr{)},$ (S11)
we replace $\sum_{\nu}A_{\mu\nu}\phi_{\nu}\to(q+qh^{2}\ell^{2}\square)\phi_{\mu}$ [11] and find
$S(\{\phi_{\mu}\})$ $=\sum_{\mu}\Bigl{(}\frac{1}{2}\phi_{\mu}(-qh^{2}\ell^{2}\square-q+\hat{m}^{2})\phi_{\mu}+\frac{u}{3!}\phi_{\mu}^{3}\Bigr{)}$
$=\sum_{\mu}\Bigl{(}qh^{2}\ell^{2}\frac{1}{2}\phi_{\mu}(-\square+m^{2})\phi_{\mu}+\frac{u}{3!}\phi_{\mu}^{3}\Bigr{)}.$ (S12)
[…]